OpenStudy (anonymous):

solve by substitution method y=5-x 2x+3y=12

OpenStudy (sgadi):

2x+3(5-x)=12 2x+15-3x=12 x=3 y=2

OpenStudy (anonymous):

can you help me i need to solve more???? please

OpenStudy (anonymous):

solve by using the substitution method x+y=6 x+4y=3